# Neural variational inference for cutting feedback during uncertainty propagation

### v1

@@ 4.1 Uniform Kullback-Leibler Approximation Rates of Conditional Normalizing Flows
We first provide results on universal approximation rates (in Kullback-Leibler divergence) of conditional normalizing flows in terms of the richness of the neural network architecture and complexity of the class of conditional densities it is approximating. These results are then used to prove the convergence of NeVI-Cut, but are also of independent importance. Much of the existing theory on fixed target approximation qualities of conditional normalizing flows is about their expressive powers (Huang et al.,, 2018; Papamakarios et al.,, 2017; Battaglia and Nicholls,, 2024) showing that the class of (conditional) normalizing flows can approximate any true (conditional) distribution in terms of weak convergence. When used as variational families, one needs to study properties of the variational solution among the conditional flow class. As the solution is obtained by minimizing the KLD, we need to quantify uniform rates of approximation of flows to a target class of densities in the KLD. This, to our knowledge, has not been developed.
We develop a theory for the setting where the conditioning variable $\eta$ is multivariate, and the observation variable $\theta$ is univariate. This is because autoregressive flows typically specify a sequence of univariate conditionals (i.e., where the observation variable is univariate while the conditioning variable is multivariate). So this setting helps to illustrate the main proof ideas and insights from the theory while balancing the technical details. We first define a suitable class of conditional densities.
Definition 1.
Let $K_{\eta}\in\mathbb{R}^{d_{\eta}}$ be some compact interval, and $K_{M}=[-M,M]\times K_{\eta}$. Define a class $\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$ of positive, continuous probability densities $p(\theta\,|\,\eta)$ supported on $\mathbb{R}\times K_{\eta}$, characterized by the following properties:
@@
    Spikiness rate (log-Lipschitz in $\theta$): There is a function $L_{S}:\mathbb{N}\to(0,\infty)$ such that,
@@
    Conditionals perturbation rate (log-Lipschitz in $\eta$): There is a function $L_{CP}:\mathbb{N}\to(0,\infty)$ such that,
@@
    Quantile growth rate: For each $p\in\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$ and $\eta\in K_{\eta}$, let
@@
    with $\Phi$ the standard normal CDF. So $F_{p}(\theta\,|\,\eta)$ is the conditional cdf, so $F^{-1}_{p}(u\,|\,\eta)$ for $u\in[0,1]$ is the conditional quantile map (with respect to the default Unif[0,1][0,1] base) and $T_{p}$ is the conditional quantile map transformed with respect to a standard Normal base. Then,
@@
      Median bound: There exists $M_{0}\geq 0$ such that $\sup_{\eta\in K_{\eta}}T(\eta,0)=M_{0}<\infty$,
@@
Assumptions $(i)$ and $(ii)$ specify that the log conditional density is local Lipschitz on compacts, with respect to both $\theta$ and $\eta$. The log-Lipschitz moduli $L_{S}(M)$ and $L_{CP}(M)$ are nonnegative and non-decreasing in $M$. The modulus $L_{S}(M)$ controls how spiky the conditional density is with respect to the observation variable $\theta$. The parameter $L_{CP}(M)$ controls how smoothly or abruptly the conditional law varies with the conditioner $\eta$. In Assumption $(iii)$, the parameters $M_{0}$, $A_{d}$, $B_{d}$ govern the global spread of the distribution via controlling the conditional quantile map. From the median and derivative bound, we immediately have the following bound on the conditional quantile map:

### v2

@@ 4.1 Uniform Kullback-Leibler Approximation Rates of Conditional Normalizing Flows
We first provide results on universal approximation rates (in Kullback-Leibler divergence) of conditional normalizing flows in terms of the richness of the neural network architecture and complexity of the class of conditional densities it is approximating. These results are then used to prove the convergence of NeVI-Cut, but are also of independent importance. Much of the existing theory on fixed target approximation qualities of conditional normalizing flows is about their expressive powers (Huang et al.,, 2018; Papamakarios et al.,, 2017; Battaglia and Nicholls,, 2024) showing that the class of (conditional) normalizing flows can approximate any true (conditional) distribution in terms of weak convergence. When used as variational families, one needs to study properties of the variational solution among the conditional flow class. Weak convergence results does not help here as the variational solution is obtained by minimizing the KLD. We need to quantify uniform rates of approximation of flows to a target class of densities in the KLD. This, to our knowledge, has not been developed.
We develop a theory for the setting where the conditioning variable $\eta$ is multivariate, and the observation variable $\theta$ is univariate. This is because autoregressive flows typically specify a sequence of univariate conditionals (i.e., where the observation variable is univariate while the conditioning variable is multivariate). So this setting helps to illustrate the main proof ideas and insights from the theory while balancing the technical details. We first define a suitable class of conditional densities. The extension to multivariate $\theta$ should be conceptually straightforward but cumbersome in terms of notation.
We first specify the true class of target conditional distributions.
Assumption 1 (True distribution class).
Let $K_{\eta}\in\mathbb{R}^{d_{\eta}}$ be some compact interval, and $K_{M}=[-M,M]\times K_{\eta}$. Let the target conditional posterior $p(\theta\,|\,\eta)$ be in a class $\mathcal{P}\left(L_{S}\left(\cdot\right),L_{CP}\left(\cdot\right),M_{0},A_{d},B_{d},c,\alpha\right)$ of positive, continuous probability densities $p(\theta\,|\,\eta)$ supported on $\mathbb{R}\times K_{\eta}$, characterized by the following properties:
@@
    Spikiness rate (log-Lipschitz in $\theta$): There is a positive increasing function $L_{S}:\mathbb{N}\to(0,\infty)$ such that,
@@
    Conditionals perturbation rate (log-Lipschitz in $\eta$): There is a positive increasing function $L_{CP}:\mathbb{N}\to(0,\infty)$ such that,
@@
    Quantile growth rate: For each $p\in\mathcal{P}\left(L_{S}\left(\cdot\right),L_{CP}\left(\cdot\right),M_{0},A_{d},B_{d},c,\alpha\right)$ and $\eta\in K_{\eta}$, let
@@
    with $\Phi$ the $N(0,1)$ CDF. So $F_{p}(\theta\,|\,\eta)$ is the conditional cdf, $F^{-1}_{p}(u\,|\,\eta)$ for $u\in[0,1]$ is the conditional quantile map (with respect to the default Unif[0,1][0,1] base) and $T_{p}$ is the conditional quantile map transformed with respect to $N(0,1)$ base. Then,
@@
      Median bound: There exists $M_{0}\geq 0$ such that $\sup_{\eta\in K_{\eta}}\left|T(\eta,0)\right|=M_{0}<\infty$,
@@
Instead of studying the properties of our method (and in general of variational conditional flows) for a specific density $p(\theta\,|\,\eta)$, we assume $p(\theta\,|\,\eta)$ satisfy Assumption 1. This enables studying the entire class of conditional families $\mathcal{P}\left(L_{S}\left(\cdot\right),L_{CP}\left(\cdot\right),M_{0},A_{d},B_{d},c,\alpha\right)$ characterized by the spikiness, conditional perturbation rate, quantile growth, and thinness bound. This is much like how Hölder or Sobolev classes provide functional envelopes for smoothness and decay and are used to derive the theory of functional estimation for these classes rather than individual functions.
Assumptions 1(i)(i) and $(ii)$ specify that the log conditional density is local Lipschitz on compacts, with respect to both $\theta$ and $\eta$. The log-Lipschitz moduli $L_{S}(M)$ and $L_{CP}(M)$ will, of course, be nonnegative and non-decreasing in $M$. The modulus $L_{S}(M)$ controls how spiky the conditional density is with respect to the observation variable $\theta$. The parameter $L_{CP}(M)$ controls how smoothly the conditional law varies with the conditioner $\eta$. Note that any locally log-Lipschitz continuous positive conditional density $p(\theta\,|\,\eta)$ supported on $\mathbb{R}\times K_{\eta}$ satisfies Assumptions 1(i)(i) and $(ii)$. The smallest $L_{S},L_{CP}$ for such a density $p(\theta\,|\,\eta)$ is induced by its own moduli and growth rates
both of which exist because of local Lipschitz continuity and positivity. So the only real assumptions are the third and fourth conditions in Assumption 1.
In Assumption $(iii)$, the parameters $M_{0}$, $A_{d}$, $B_{d}$ govern the global spread of the distribution via controlling the conditional quantile map. From the median and derivative bounds, we immediately have the following bound on the conditional quantile map: